OpenStudy (anonymous):

simplify -5a+7-2(4a-3)-6a+4(2-3a)

OpenStudy (angela210793):

-5a+7-8a+6-6a+8-12a=-31a+21 -31a=-21 31a=21 a=21/31

OpenStudy (angela210793):

only simplify?

OpenStudy (anonymous):

yes

OpenStudy (angela210793):

then =-31a+21

OpenStudy (anonymous):

Thank you